OpenStudy (anonymous):

Write the equation for the differential, written dy, for y = sec (5x).Simplify, where possible.

OpenStudy (mathmale):

1) Differentiate y=sec 5x in the usual way. 2) Multiply both sides of the resulting equation by dx.

OpenStudy (anonymous):

could you show me how?

OpenStudy (anonymous):

\[\frac{ dy }{ dx } = \frac{ d }{ dx }(\sec5x)\] \[\sec5x*\tan5x*\frac{ d }{ dx }(5x)\] hope u can do it now

OpenStudy (anonymous):

would the answer be dy=5y tan(5x)dx?

OpenStudy (anonymous):

yeah